OpenStudy (anonymous):

radicals: (√5+√2)^2?

myininaya (myininaya):

who will win? jimmy or jim?

myininaya (myininaya):

jim vs jim!

OpenStudy (lalaly):

lol myininaya hehe

jimthompson5910 (jim_thompson5910):

\[\large (\sqrt{5}+\sqrt{2})^2\] \[\large (\sqrt{5})^2+2\sqrt{5}\sqrt{2}+(\sqrt{2})^2\] \[\large 5+2\sqrt{10}+2\] \[\large 7+2\sqrt{10}\] So \[\large (\sqrt{5}+\sqrt{2})^2=7+2\sqrt{10}\]

myininaya (myininaya):

looks good jim

OpenStudy (anonymous):

why would it be 7+2, instead of 9?

OpenStudy (anonymous):

= 5 + √10 + √10 + 2 = 7 + 2√10

myininaya (myininaya):

can you combine terms here: 7+2x?

myininaya (myininaya):

say no

OpenStudy (anonymous):

i'm slow tonight

OpenStudy (anonymous):

ah i see. thank you!

myininaya (myininaya):

7+2apples is just 7 +2apples

myininaya (myininaya):

jimmy you are still a winner

OpenStudy (anonymous):

ah - thank you myininaya

OpenStudy (anonymous):

ts my grandaughters birthday today so i celebrated a little

myininaya (myininaya):

eat some cake for me!